OpenStudy (anonymous):

An exam is given to 100 students. The population of the class was 40 male students and 60 female students. Fifty-six students passed the exam. Of the students that pass the class 20 of them were men, and 36 were women. In order to get feedback on the course, a student is selected to give an interview about the exam. Find the probability that the student passed, given that the student is male. Since more women passed the exam than men, is it more likely that a student will pass the exam if they are female?

OpenStudy (anonymous):

Anybody got any idea?

jimthompson5910 (jim_thompson5910):

how many males are there total?

OpenStudy (anonymous):

40

jimthompson5910 (jim_thompson5910):

how many of those males passed

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

so if we're told that you picked a male, then 20/40 = 1/2 is the probability he passed

jimthompson5910 (jim_thompson5910):

in probability notation, it would be P(passed|male) = 1/2

OpenStudy (anonymous):

Just didn't look that simple

jimthompson5910 (jim_thompson5910):

yeah it's not too bad

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

second part is correct? Female is more likely to pass?

jimthompson5910 (jim_thompson5910):

yes because P(pass|female) = 36/60 = 3/5 = 0.6 which is greater than 1/2 or 0.5